OpenStudy (miss.rose):

54 Miss.Rose Medals 0 The first question is: Solve. −s3 ≥6 s ≥ –2 s ≤ –2 s ≥ –18 s ≤ –18 This one? Question 2: Identify the solution set of the inequality, using the given replacement set. x < –4; {–10, –4.3, –4, –3.9, 2, 6.5} {–10, –4.3, –4} {–4, –3.9, 2} {–10, –4.3} This one? {–3.9, 2} Question 3: Find the inequality whose solution is the following: -10-5 -40 y + 3 ≥ –1 y + 3 > –1 y + 3 ≤ –1 y + 3 < –1 This one? Question 4: Find the inequality whose solution is the following: -10-50510-7 y + 5 –2 y + 5 ≥ –2 y + 5

OpenStudy (miss.rose):

@zoeeh127

OpenStudy (anonymous):

2 is So, you have an inequality: x<−4 You also have a replacement set: {–10,–4.3,–4,–3.9,2,6.5} The solution set is just all of the numbers in the replacement set that make the inequality true. -10 < -4, that one works. -4.3 < -4, that one works. -4 < -4, nope, that doesn't work. -3.9, 2, 6.5 all fail as well. So the solution set is simply {−10,−4.3}

OpenStudy (anonymous):

sorry that's the only one I can get

OpenStudy (miss.rose):

It's okay. Do you know anyone else who can help? @tuxadog12

OpenStudy (anonymous):

UM one sec let me see

OpenStudy (anonymous):

I just asked someone I just have to see if they will come

OpenStudy (miss.rose):

Okay.

OpenStudy (miss.rose):

These are the questions I need help with.

OpenStudy (anonymous):

sorry im only in 7th grade i dont understand these

OpenStudy (anonymous):

ok thanks for trying

OpenStudy (miss.rose):

I am in the 7th grade too. So I don't know how you don't understand them.

OpenStudy (anonymous):

sorry

OpenStudy (anonymous):

I couldn't be much more help

OpenStudy (miss.rose):

It's okay. I have someone who can try. @JoeDeWise Can you help?

OpenStudy (anonymous):

ok gtg

OpenStudy (anonymous):

well i guess i haven't learned it yet

OpenStudy (anonymous):

No lo sé

OpenStudy (anonymous):

I do not know

OpenStudy (miss.rose):

@rebeccaxhawaii

OpenStudy (anonymous):

@Mehek14

rebeccaxhawaii (rebeccaxhawaii):

im sorry i havent done this in so long

OpenStudy (miss.rose):

It's okay. Do you know anyone who can?

OpenStudy (anonymous):

Question 1: \[−s3 ≥6 \] divide both sides by -3, and since the divisor is negative, we have to flip the operator sign: \[s \le \frac{ 6 }{ -3 }\] and we'll get: \[s \le -2\] which happens to be B

OpenStudy (anonymous):

For Question 2, you're right @Miss.Rose

OpenStudy (anonymous):

For Question 3, we have to suppose an example equation first: \[-10, -5, -40\] \[y + 3 = -1\] \[y = -4\] now we compare each given value's relation to y = -4 \[-10 < -4\] \[-5 < -4\] \[-40 < -4\] we see that all three values are less than -y, so we can conclude: \[y + 3 < -1\] which happens to be D as you suggested, but keep in mind that C is also a possibility.